• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Ibrahim

    Log entry time 04:56:08 on June13,2002

    Entry number 85970

    This entry is a followup to: 85960

    keyword=Target Problem


    At about 1:10 a.m. I noticed that all the charts are flatlined, so I decided to reboot the IOC with beam on since it was stable for a considerable time. After the IOC came back we noticed that the temperature of the LD2 loop is falling quikly so I started to use the Aux. Heater to control it and called the MCC to take the beam off untill I check why the PID is not functional. I checked all the different seetings that I knew at that time and everything seemed normal to me except the PID. Thus I decided to reboot the IOC again as a final trial before I call K. McCormick (the target-on-call). This time also I did not see much difference, so I called K. McCormic and she told me to check several things and put by hand all the PID parameters. After more than one hour of continous adjustments we finally saw some indiction of stablity. In the end we got the temperatures of both loops as set. Only one problem we could not fix which is the readings of the following LD2 temperature sensors are not correct on the alarm set point window:

    ITC502_5_Sensor_1_R
    ITC502_5_Sensor_2_R
    ITC502_6_Sensor_1_R
    ITC502_6_Sensor_2_R

    although the target is currently controlled by them (as i can see loop1 is at 22.00K).
    It might be a problem of the GUI which needs to be rebooted. Anyway we are taking data right now untill some expert check this problem.



    FIGURE 1